A 200 mL solution of sugar water has a sugar concentration of 40%. How many mL of pure water must be added to the solution to reduce the concentration to 10%?
Answer by mananth(16949)

-------------- percent ---------------- Amount
Sugar solution 40 ---------------- 200 ml
water 0 ---------------- x ml
Total 10 ---------------- 200 + x ml

40 * 200 + 0 x = 10 ( 200 + x )
8000 + 0 x = 2000 + 10 x
0 x -10 x = -8000 + 2000
-10 x = -6000
/ -10
x = 600 ml water

m.ananth@hotmail.ca